The following data represent a random sample of 9 marks (out of 10) on a statistics quiz. The marks are normally distributed with a standard deviation of 2. Estimate the population mean with 90% confidence.

7 9 7 5 4 8 3 10 9


1
Expert's answer
2021-11-26T12:29:26-0500

Mean "\bar x=\frac{7+9+7+5+4+8+3+10+9}{9}=6.889."

"90\%CI=(\bar x-z_{0.05}\frac{\sigma}{\sqrt{n}},\bar x+z_{0.05}\frac{\sigma}{\sqrt{n}})="

"=(6.889-1.645\frac{2}{\sqrt{9}},6.889+1.645\frac{2}{\sqrt{9}})="

"=(5.792,7.986)."